OpenStudy (anonymous):

ahh, okay word problems+me=mess Help me with my last homework question please! Use the formula h = -16 t2 + 48t + 160 to determine when the ball's height is 192 feet. The formula describes the ball's height above the ground, h, in feet, t seconds after you throw the ball straight up from a rooftop 160 feet high.

OpenStudy (goformit100):

use pythagoras theorem

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

Im stuck on it but I will try again

OpenStudy (anonymous):

I am trying to factor 192=16t^2+48t+160 correct?

OpenStudy (phi):

first subtract 192 from both sides

OpenStudy (phi):

to get -16t^2+48t-32=0 now simplify by dividing both sides by -16 (all terms!) to get t^2-3t+2=0 (I just noticed you left the minus sign of the -16t^2 term)

OpenStudy (phi):

you follow?

OpenStudy (anonymous):

I did and I figured it out! thank you so much t=1

OpenStudy (phi):

yes t^2-3t+2 factors into (t-1)(t-2)= 0 t= 1 or t=2 t=1 is on the way up t=2 is when the ball reaches 192 ft on the way down.

OpenStudy (phi):

so you were on the right track with I am trying to factor 192=16t^2+48t+160 correct? except you should have -16t^2. to solve, make it a quadratic equal to 0 (by subtracting 192 from both sides), then simplify by dividing everything by -16. After that, factor it to get the answer.